OpenStudy (kimo_indie):

What is the correct classification of the system of equations below? 2x + 6y = 8 y - 2x = 2 parallel coincident intersecting

OpenStudy (liv1234):

What do you think the answer is?

OpenStudy (alexandervonhumboldt2):

parallel lines have same slope Coincident. Two lines or shapes that lie exactly on top of each other.

OpenStudy (alexandervonhumboldt2):

2x + 6y = 8 => 6y=8-2x => y=-1/3x+4/3 y - 2x = 2 => y=2+2x => y=2x+2

OpenStudy (alexandervonhumboldt2):

these lines are just intersecting

OpenStudy (kimo_indie):

thnx for explaing @AlexandervonHumboldt2